Even payments of $600 are made in year 0, 1, 2, 3, 4 and 5 into an account that makes 10% interest. Instead of that investment scheme, find the amount of two equal payments made in year 1 and year 4 that would give an equivalent balance.
Answer #1

set the equation as below

600+600/(1+10%)^1+600/(1+10%)^2+600/(1+10%)^3+600/(1+10%)^4+600/(1+10%)^5=A/(1+10%)^1+A/(1+10%)^4

600+600/(1+10%)^1+600/(1+10%)^2+600/(1+10%)^3+600/(1+10%)^4+600/(1+10%)^5=A*(1/(1+10%)^1+1/(1+10%)^4)

A=(600+600/(1+10%)^1+600/(1+10%)^2+600/(1+10%)^3+600/(1+10%)^4+600/(1+10%)^5)/((1/(1+10%)^1+1/(1+10%)^4))

=1805.45 (answer)